EXPERIMENTAL DETECTION AND THEORETICAL CHARACTERIZATION OF THE H_{2}-NH(X) VAN DER WAALS COMPLEX

W. M. FAWZY, G. KERENSKAYA, and \underline{M. C. HEAVEN}, Department of Chemistry and Cherry L. Emerson Center for Scientific Computation, Emory University, Atlanta, GA 30322.

The H_{2}-NH(X) van der Waals complex has been examined using ab initio theory and detected via fluorescence excitation spectroscopy of the A^{3}Π - X^{3}Σ^{−} transition. Electronic structure calculations show that the minimum energy geometry corresponds to collinear H_{2}-NH(X), with a well depth of D_{e}=116 cm^{−1}. The potential energy surface supports a secondary minimum for a T-shaped geometry, where the H atom of NH points towards the middle of the H_{2} bond (C_{2}v point group). For this geometry the well depth is 73 cm^{−1}. Laser excitation spectra for the complex show transitions to the H_{2}+NH(A) dissociative continuum. The onset of the continuum establishes a binding energy of D_{0}=32±2 cm^{−1} for H_{2}-NH(X). Fluorescence from bound levels of H_{2}-NH(A) was not detected, most probably due to rapid reactive decay (H_{2}-NH(A) → H+NH_{2}). The complex appears to be a promising candidate for studies of the photo-initiated H_{2}+NH abstraction reaction under conditions were the reactants are pre-aligned by the van der Waals forces.